# Sharp concentration inequality for the sum of random variables

### v1

@@ 1 Maximally nonincreasing operator framework and main results
2 Mixed slot variables to sharpnen sum concentration tail
Given $n\in\mathbb{N}$, $p\in]0,1[$ and a probability measure $\mu$ on $\mathbb{R}$, introduce the following notations to prepare the construction [^10] The choice of $u_{0}=u_{0}^{\prime}$ is not important because $\mu(]\inf T_{\mu}(p),\sup T_{\mu}(p)[)=0$.. Figure 2 gives a graphical representation of this choice:
  • $p_{0}=p_{0}^{\prime}:=p$ and, for all $i\in[n]$,
    $p_{i}:=\frac{n-i}{n}p,\qquad p_{i}^{\prime}:=\frac{i}{n}+\frac{n-i}{n}p,$
  • for all $i\in[n]$, pick any
    $u_{i}\in T_{X}(p_{i})$ $\operatorname{and}$ $u_{i}^{\prime}\in T_{X}(p_{i}^{\prime}),$
    if $T_{X}(1)=\emptyset$ then take $u_{n}^{\prime}=$-$\infty$, and if $T_{X}(0)=\emptyset$, take $u_{n}=+\infty$.
  • pick $u_{0}=u_{0}^{\prime}\in T_{X}(p)$,
Figure 2: (Left) detail of the choice of of the different parameters for a given tail quantile operator and (Right) cyclic mixing of slot variables: value of $X_{1},\ldots,X_{n},\sum_{k=1}^{n}X_{k}$ depending on the values of $\epsilon_{1},\ldots,\epsilon_{n},\epsilon_{1}^{\prime},\ldots,\epsilon_{n}^{\prime}$.
Note that:
  • (P1a)
    $\infty\leq u_{n}^{\prime}\leq u_{n-1}^{\prime}\leq\dots\leq u_{0}^{\prime}=u_{0}\leq\dots\leq u_{n-1}\leq u_{n}\leq+\infty$,
  • (P1b)
    The set $\bigcup_{i\in[n]}[u_{i}^{\prime},u_{i-1}^{\prime}[\ \cup\ \bigcup_{i\in[n]}[u_{i-1},u_{i}[$ forms a partition of $\mathbb{R}$,
For any $i\in[n]$, we construct two probability measures $\mu_{i},\mu_{i}^{\prime}$ on $\mathbb{R}$ through their survival function, respectively by setting, forall $t\in\mathbb{R}$:
$S_{\mu_{i}}(t)=\begin{cases}1&\text{if }t<u_{i-1},\\ \frac{S_{\mu}(t)-p_{i}}{p_{i-1}-p_{i}}&\text{if }u_{i-1}<t<u_{i},\\ 0&\text{if }t>u_{i}\end{cases}$ $\operatorname{and}$ $S_{\mu_{i}^{\prime}}(t)=\begin{cases}1&\text{if }t<u_{i}^{\prime},\\ \frac{S_{\mu}(t)-p_{i-1}^{\prime}}{p_{i}^{\prime}-p_{i-1}^{\prime\prime}}&\text{if }u_{i}^{\prime}<t<u_{i-1}^{\prime},\\ 0&\text{if }t>u_{i-1}^{\prime},\end{cases}$ (8)
where the intermediate set value denotes, in both cases, an affine image of the interval $S_{\mu}(t)$. One can check that with this definitions:
  • (P2a)
    if $A\cap[u_{i-1},u_{i}]=\emptyset$ (resp. $A\cap[u_{i}^{\prime},u_{i-1}^{\prime}]=\emptyset$), then $\mu_{i}(A)=0$ (resp. $\mu_{i}^{\prime}(A)=0$),
  • (P2b)
    if $u_{i}=u_{i-1}$ (resp. $u_{i}^{\prime}=u_{i-1}^{\prime}$), then $\mu_{i}(A)=\mathbbm{1}(u_{i}\in A)$ (resp. $\mu_{i}^{\prime}(A)=\mathbbm{1}(u_{i}^{\prime}\in A)$).
As defined, $\mu_{i}$ and $\mu_{i}^{\prime}$ can be viewed as “slot restrictions” of $\mu$, as stated by the next lemma.
Lemma 2.
Given $i\in[n]$ and $s\in]0,1[$,
$T_{\mu_{i}}(s)=T_{\mu}\!\left(p_{i}+\frac{p}{n}s\right),\qquad T_{\mu_{i}^{\prime}}(s)=T_{\mu}\!\left(p_{i-1}^{\prime}+\frac{1-p}{n}s\right).$
Proof.
We prove the first identity (the second one is analogous). If $u_{i}=u_{i-1}$, then by (P2b) the measure $\mu_{i}$ is the Dirac mass at $u_{i}$, hence $T_{\mu_{i}}(s)=\{u_{i}\}$ for all $s\in]0,1[$. Since $T_{\mu}$ is constant on $(p_{i},p_{i-1})$ in that case, the identity follows.
Assume now that $u_{i}\neq u_{i-1}$. Following the formulas given in (8), for any $s\in]0,1[$ and $t\in\mathbb{R}$,
$s\in S_{\mu_{i}}(t)\quad\Longleftrightarrow\quad p_{i}+(p_{i-1}-p_{i})s\in S_{\mu}(t)\quad\Longleftrightarrow\quad t\in T_{\mu}\!\left(p_{i}+\frac{p}{n}s\right),$
since $p_{i-1}-p_{i}=\frac{p}{n}$. This proves $T_{\mu_{i}}(s)=T_{\mu}\!\left(p_{i}+\frac{p}{n}s\right)$. ∎
Define now “independent slot variables” $W_{n},W_{n-1},\dots,W_{1},W^{\prime}_{1},\dots,W^{\prime}_{n-1},W^{\prime}_{n}$ such that for any $i\in[n]$:
$S_{W_{i}}=S_{\mu_{i}}\qquad\operatorname{and}\qquad S_{W_{i}^{\prime}}=S_{\mu_{i}^{\prime}}.$
We call them “slot variables” since, by (P2a) :
$W_{i}\in[u_{i-1},u_{i}]$ $\operatorname{and}$ $W_{i}^{\prime}\in[u_{i}^{\prime},u_{i-1}^{\prime}]$ $a.s.$ (9)
Lemma 3.
With the notation of (6),
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}\right]=\mathcal{H}(T_{\mu})(p)\qquad\text{and}\qquad\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}\right]=\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p).$
Proof.
By Lemma 2 and the identity $\mathbb{E}[Y]=\int_{0}^{1}T_{Y}(s)\,ds$, true for integrable $Y$ (see for instance (Rockafellar and Royset, 2014, (3.4))),
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}\right]$ $=\frac{1}{n}\sum_{i=1}^{n}\int_{0}^{1}T_{\mu_{i}}(s)\,ds=\frac{1}{n}\sum_{i=1}^{n}\frac{n}{p}\int_{p_{i}}^{p_{i-1}}T_{\mu}(u)\,du=\frac{1}{p}\int_{0}^{p}T_{\mu}(u)\,du=\mathcal{H}(T_{\mu})(p).$
Similarly,
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}\right]$ $=\frac{1}{n}\sum_{i=1}^{n}\int_{0}^{1}T_{\mu_{i}^{\prime}}(s)\,ds=\frac{1}{n}\sum_{i=1}^{n}\frac{n}{1-p}\int_{p_{i-1}^{\prime}}^{p_{i}^{\prime}}T_{\mu}(u)\,du$
$=\frac{1}{1-p}\int_{p}^{1}T_{\mu}(u)\,du=\frac{\mathbb{E}[\mu]}{1-p}-\frac{p}{1-p}\mathcal{H}(T_{\mu})(p)=\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p).$
where we recall the notation introduced in (6): $\Delta_{\mu}(p):=\frac{\mathcal{H}(T_{\mu})(p)-\mathbb{E}[X]}{1-p}$. ∎
Actually, we have a law of large number (LLN) for those slot variables
Lemma 4.
If we assume that $\mu$ admits an expectation than, for any $p\in]0,1[$ and any $t>0$:
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}-\mathcal{H}(T_{\mu})(p)\right|\geq t\right)\underset{n\to\infty}{\longrightarrow}0$
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}-\mathcal{H}(T_{\mu})(p)+\Delta_{\mu}(p)\right|\geq t\right)\underset{n\to\infty}{\longrightarrow}0$
Proof.
We just prove the first result since the LLN for $\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}$ is proven the same way. Let us first exclude the last elements of the sum that can be unbounded, then we can bound alsmost surely thanks to (9):
$\frac{1}{n}\sum_{i=1}^{n-1}u_{i}\leq\frac{1}{n}\sum_{i=1}^{n-1}W_{i}\leq\frac{1}{n}\sum_{i=1}^{n-1}u_{i-1},$
and by definition of the integral of operators (which is a natural extention of the inntegral of measurable mappings defined as the limit of integral of simple functions as we see them apppear here), we see that:
$\frac{1}{n}\sum_{i=1}^{n-1}W_{i}\underset{n\to\infty}{\longrightarrow}0\qquad a.s$
Besides we know from definition of $S_{\mu_{n}}$ in (8) that ($p_{n}=0$):
$\mathbb{P}\left(\left|\frac{W_{n}}{n}\right|\geq t\right)=S_{\mu_{i}}(tn)\leq\frac{S_{\mu}(tn)}{p_{n-1}}=\frac{n}{p}S_{\mu}(tn)=\frac{n}{p}\int_{tn}^{\infty}d\mu(u)\leq\frac{1}{pt}\int_{tn}^{\infty}ud\mu(u)\underset{n\to\infty}{\longrightarrow}0,$
by continuity of the integral since $u\to u$ is integrable under $\mu$ by assumption. ∎
We now construct $n$ dependent random variables following law $\mu$ by mixing the slot variables. Let $(\epsilon_{1},\ldots,\epsilon_{n},\epsilon_{1}^{\prime},\ldots,\epsilon_{n}^{\prime})\in\{0,1\}^{2n}$ be a random vector, independent of $(W_{i})_{i\in[n]}$ and $(W_{i}^{\prime})_{i\in[n]}$, such that:
  1. 1.
    $\mathbb{P}\!\left(\sum_{i=1}^{n}\epsilon_{i}+\sum_{i=1}^{n}\epsilon_{i}^{\prime}=1\right)=1$,
  2. 2.
    $\mathbb{P}(\epsilon_{i}=1)=\frac{p}{n}$ and $\mathbb{P}(\epsilon_{i}^{\prime}=1)=\frac{1-p}{n}$ for all $i\in[n]$.
For any $k,i\in[n]$, define the cyclic permutation
$\sigma_{k}(i)=\begin{cases}i+k-1,&\text{if }i+k-1\leq n,\\ i+k-n-1,&\text{if }n+1\leq i+k-1\ (\leq 2n-1).\end{cases}$
Finally, following the description given on Figure 2, define
$\forall k\in[n]:\qquad X_{k}=\sum_{i=1}^{n}\epsilon_{\sigma_{k}(i)}W_{i}+\sum_{i=1}^{n}\epsilon_{\sigma_{k}(i)}^{\prime}W_{i}^{\prime}.$ (10)
Lemma 5.
For all $k\in[n]$, $X_{k}\sim\mu$.
Proof.
Since $(\epsilon_{i})_{i}$ and $(\epsilon_{i}^{\prime})_{i}$, $(W_{i})_{i}$ and $(W_{i}^{\prime})_{i}$ are all independent, we can express, for any $t\in\mathbb{R}$:
$S_{X_{k}}(t)=\sum_{i=1}^{n}\mathbb{P}(\epsilon_{\sigma_{k}(i)}=1)\,S_{\mu_{i}}(t)+\sum_{i=1}^{n}\mathbb{P}(\epsilon_{\sigma_{k}(i)}^{\prime}=1)\,S_{\mu_{i}^{\prime}}(t)$
Now if, say, $t\in[u_{i-1},u_{i}]$, we know that
  • $\forall j<i$: $S_{\mu_{j}}(t)=0$,
  • $\forall j\in[n]$: $S_{\mu_{j}^{\prime}}(t)=0$ and $P(\epsilon_{\sigma_{k}(i)})=1$,
  • $\forall j>i$: $S_{\mu_{j}}(t)=1$,
  • $S_{\mu_{i}}(t)=\frac{S_{\mu}(t)-p_{i}}{p_{i-1}-p_{i}}$.
That allows us to evaluate:
$S_{X_{k}}(t)=\frac{p}{n}\frac{S_{\mu}(t)-p_{i}}{p_{i-1}-p_{i}}+\frac{n-i}{n}\frac{p}{n}=S_{\mu}(t),$
thanks to the identity $p_{i-1}-p_{i}=\frac{p}{n}$ and the definition of $p_{i}$. The same thing happens if $t$ belongs to any other slot given in (P1b) . ∎
Proof of Theorem 2.
Using (10) and the fact that $\sum_{k=1}^{n}\epsilon_{\sigma_{k}(i)}=\sum_{j=1}^{n}\epsilon_{j}$ for each fixed $i$, we obtain
$\frac{1}{n}\sum_{k=1}^{n}X_{k}$ $=\frac{1}{n}\left(\sum_{i=1}^{n}\left(\sum_{k=1}^{n}\epsilon_{\sigma_{k}(i)}\right)W_{i}+\sum_{i=1}^{n}\left(\sum_{k=1}^{n}\epsilon^{\prime}_{\sigma_{k}(i)}\right)W^{\prime}_{i}\right)$
$=\epsilon\left(\frac{1}{n}\sum_{i=1}^{n}W_{i}\right)+(1-\epsilon)\left(\frac{1}{n}\sum_{i=1}^{n}W^{\prime}_{i}\right)$ (11)
where $\epsilon:=\sum_{j=1}^{n}\epsilon_{j}\in\{0,1\}$ satisfies $\mathbb{P}(\epsilon=1)=p$ and $\mathbb{P}(\epsilon=0)=1-p$.
Using (2), for any $t$ such that
$\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p)<t<\mathcal{H}(T_{\mu})(p),$
we have with Lemma 4:
$\mathbb{P}\left(\frac{1}{n}\sum_{k=1}^{n}X_{k}\geq t\right)$ $=p\,\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}W_{i}\geq t\right)+(1-p)\,\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}\geq t\right)\underset{n\to\infty}{\longrightarrow}p.$
Similarly,
$\forall t<\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p):\ \ \mathbb{P}(Y_{n}\geq t)\to 1\qquad\operatorname{and}\qquad\forall t>\mathcal{H}(T_{\mu})(p):\ \ \mathbb{P}(Y_{n}\geq t)\to 0,$
which is exactly the profile of $S_{\mu,p}$. ∎

### v2

@@ 1 Introduction and main results
2 Sharpness for triangular array of marginals
In this section, we show that, under a uniform envelope assumption on marginal tails, the universal concentration bound from Theorem 1.2, applied for each $n$ to a triangular array of random variables $(X^{(n)}_{i})_{i\in[n]}$, is asymptotically sharp at every level $p\in]0,1[$. We will denote for simplicity
$\bar{X}^{(n)}=\frac{1}{n}\sum_{i=1}^{n}X_{i}^{(n)}.$
Moreover, given a family of $n$ measures $\nu=(\nu_{1},\ldots,\nu_{n})$, we denote the averaged tail-quantile operator
$T_{\nu}:=\frac{1}{n}\sum_{i=1}^{n}T_{\nu_{i}}\in\mathcal{M}_{\downarrow},$
and set
$b_{\nu}(p):=\mathcal{H}(T_{\nu})(p),\qquad\Delta_{\nu}(p):=\frac{b_{\nu}(p)-\int_{0}^{1}T_{\nu}(u)\,du}{1-p},\qquad a_{\nu}(p):=b_{\nu}(p)-\Delta_{\nu}(p).$
Theorem 2.1 (Asymptotic sharpness for non-identically distributed marginals).
Fix $p\in]0,1[$. For each $n\in\mathbb{N}$, let $\mu^{(n)}=(\mu^{(n)}_{1},\ldots,\mu^{(n)}_{n})$ be a family of probability measures on $\mathbb{R}$ such that there exists a measurable function $h:]0,1[\to\mathbb{R}_{+}$ such that
$\int_{0}^{1}h(u)\,du<\infty\qquad\text{and}\qquad\forall n\in\mathbb{N},\ \forall i\in[n],\ \forall u\in]0,1[:\quad T_{\mu_{i}^{(n)}}(u)\subset[-h(u),h(u)].$ (8)
Then, for each $n$, there exists a family $(X^{(n)}_{1},\ldots,X^{(n)}_{n})$ such that $X^{(n)}_{i}\sim\mu^{(n)}_{i}$ for all $i\in[n]$, and for any $\eta>0$ and any sequence $(t_{n})_{n\in\mathbb{N}}$ such that for all large $n$,
$t_{n}\leq a_{\mu^{(n)}}(p)-\eta,$ or $a_{\mu^{(n)}}(p)+\eta\leq t_{n}\leq b_{\mu^{(n)}}(p)-\eta,$ or $t_{n}\geq b_{\mu^{(n)}}(p)+\eta,$
one has respectively
$\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 1,$ $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to p,$ $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 0.$
Assumption (8) allows to use the dominated convergence theorem to set a law of large numbers in the proof of the theorem, it also ensures that $X^{(n)}_{i}$ for $n\in\mathbb{N}$, $i\in[n]$ all have finite expectations.
Proof 2.2.
Step 1: Quantile representation (push-forward form). For each $n\in\mathbb{N}$ and each $i\in[n]$, choose a measurable selection $\phi_{i}^{(n)}:[0,1]\to\mathbb{R}$ such that $\phi_{i}^{(n)}(u)\in T_{\mu_{i}^{(n)}}(u)$ for all $u\in[0,1]$. Equivalently, $\mu_{i}^{(n)}$ is the push-forward of the Lebesgue measure on $[0,1]$ through $\phi_{i}^{(n)}$.
Let $\epsilon$ be a Bernoulli random variable with $\mathbb{P}(\epsilon=1)=p$, and let $(U_{i})_{i\in[n]}$ be i.i.d. $\mathrm{Unif}(0,1)$, independent of $\epsilon$. Define, for all $i\in[n]$,
$U_{i}^{\epsilon}:=\epsilon\,(pU_{i})+(1-\epsilon)\,\bigl(p+(1-p)U_{i}\bigr),\qquad X_{i}^{(n)}:=\phi_{i}^{(n)}(U_{i}^{\epsilon}).$ (9)
By construction, conditional on $\epsilon$, the variables $(U_{i}^{\epsilon})_{i\in[n]}$ are i.i.d. and uniform on $[0,p]$ if $\epsilon=1$ and uniform on $[p,1]$ if $\epsilon=0$; moreover, unconditionally, each $U_{i}^{\epsilon}$ is $\mathrm{Unif}(0,1)$. Therefore, for each $i\in[n]$, $X_{i}^{(n)}\sim\mu_{i}^{(n)}$.
Step 2: Conditional expectations and identification of $a_{\mu^{(n)}}(p),b_{\mu^{(n)}}(p)$.
On $\{\epsilon=1\}$, $U_{i}^{\epsilon}=pU_{i}$, hence
$\mathbb{E}\bigl[X_{i}^{(n)}\mid\epsilon=1\bigr]=\int_{0}^{1}\phi_{i}^{(n)}(pu)\,du=\frac{1}{p}\int_{0}^{p}\phi_{i}^{(n)}(u)\,du=\mathcal{H}(T_{\mu_{i}^{(n)}})(p),$
where the last equality follows from the definition of the Aumann integral and the fact that $\mathcal{H}(T_{\mu_{i}^{(n)}})(p)$ is a singleton for $p\in]0,1[$. Averaging over $i$ and using linearity of the integral,
$\mathbb{E}\bigl[\bar{X}^{(n)}\mid\epsilon=1\bigr]$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\bigl[X_{i}^{(n)}\mid\epsilon=1\bigr]$
$=\frac{1}{p}\int_{0}^{p}\left(\frac{1}{n}\sum_{i=1}^{n}T_{\mu_{i}^{(n)}}(u)\right)\,du=\mathcal{H}\!\left(T_{\mu^{(n)}}\right)(p)=b_{\mu^{(n)}}(p).$ (10)
Similarly, on $\{\epsilon=0\}$, $U_{i}^{\epsilon}=p+(1-p)U_{i}$, hence
$\mathbb{E}\bigl[X_{i}^{(n)}\mid\epsilon=0\bigr]=\int_{0}^{1}\phi_{i}^{(n)}\bigl(p+(1-p)u\bigr)\,du=\frac{1}{1-p}\int_{p}^{1}\phi_{i}^{(n)}(u)\,du,$
and summing over $i$ gives
$\mathbb{E}\bigl[\bar{X}^{(n)}\mid\epsilon=0\bigr]=\frac{1}{1-p}\int_{p}^{1}\left(\frac{1}{n}\sum_{i=1}^{n}T_{\mu_{i}^{(n)}}(u)\right)\,du=\frac{1}{1-p}\int_{p}^{1}T_{\mu^{(n)}}(u)\,du.$ (11)
On the other hand, by the identity $\mathbb{E}[Z]=\int_{0}^{1}T_{Z}(u)\,du$ for integrable $Z$ (see [Rockafellar and Royset(2014), (3.4)]),
$\mathbb{E}\bigl[\bar{X}^{(n)}\bigr]=\frac{1}{n}\sum_{i=1}^{n}\int_{0}^{1}T_{\mu_{i}^{(n)}}(u)\,du=\int_{0}^{1}T_{\mu^{(n)}}(u)\,du.$
Combining this with (2.2) and (11) yields
$\mathbb{E}\bigl[\bar{X}^{(n)}\mid\epsilon=0\bigr]=\frac{\mathbb{E}\bigl[\bar{X}^{(n)}\bigr]-p\,b_{\mu^{(n)}}(p)}{1-p}=b_{\mu^{(n)}}(p)-\frac{b_{\mu^{(n)}}(p)-\int_{0}^{1}T_{\mu^{(n)}}(u)\,du}{1-p}=a_{\mu^{(n)}}(p).$
Step 3: Conditional law of large numbers. We show that $\bar{X}^{(n)}\mid\{\epsilon=1\}\to b_{\mu^{(n)}}(p)$ in probability; the case $\epsilon=0$ is identical and yields $\bar{X}^{(n)}\mid\{\epsilon=0\}\to a_{\mu^{(n)}}(p)$.
Fix $t>0$ and work on the event $\{\epsilon=1\}$. Then $(X_{i}^{(n)})_{i\in[n]}$ are independent. For $M>0$, define the truncations
$X_{i}^{(n,M)}:=(-M)\vee X_{i}^{(n)}\wedge M,\qquad\bar{X}^{(n,M)}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}^{(n,M)}.$
Fix $\delta>0$ and, without loss of generality, assume $\delta\leq 1$. Using a union bound, we obtain
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-b_{\mu^{(n)}}(p)\bigr|\geq t\ \bigm|\ \epsilon=1\Bigr)$ $\leq\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-\bar{X}^{(n,M)}\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)$ (12)
$\quad+\mathbb{P}\Bigl(\bigl|\bar{X}^{(n,M)}-\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)$
$\quad+\mathbb{P}\Bigl(\bigl|\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]-b_{\mu^{(n)}}(p)\bigr|\geq t/3\Bigr).$
We choose $M$ large enough so that the first term is at most $\delta/2$ (uniformly in $n$) and the third term is equal to 0; then we let $n\to\infty$ to control the second term via Hoeffding’s inequality.
We start with the bound
$\bigl|\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]-\mathbb{E}[\bar{X}^{(n)}\mid\epsilon=1]\bigr|$ $\leq\mathbb{E}\bigl[|\bar{X}^{(n)}-\bar{X}^{(n,M)}|\mid\epsilon=1\bigr]$
$\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\bigl[|X_{i}^{(n)}-X_{i}^{(n,M)}|\mid\epsilon=1\bigr],$ (13)
and $\mathbb{E}[\bar{X}^{(n)}\mid\epsilon=1]=b_{\mu^{(n)}}(p)$ by (2.2). Hence the third term in (12) is zero as soon as
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\bigl[|X_{i}^{(n)}-X_{i}^{(n,M)}|\mid\epsilon=1\bigr]<t/3.$
Moreover, since $|X_{i}^{(n)}-X_{i}^{(n,M)}|\leq|X_{i}^{(n)}|\,\mathbf{1}_{\{|X_{i}^{(n)}|>M\}}$, using the quantile representation on $\{\epsilon=1\}$ we have
$\mathbb{E}\bigl[|X_{i}^{(n)}-X_{i}^{(n,M)}|\mid\epsilon=1\bigr]\leq\frac{1}{p}\int_{0}^{p}|\phi_{i}^{(n)}(u)|\,\mathbf{1}_{\{|\phi_{i}^{(n)}(u)|>M\}}\,du\leq\frac{1}{p}\int_{0}^{p}h(u)\,\mathbf{1}_{\{h(u)>M\}}\,du.$
The right-hand side tends to 0 as $M\to\infty$ by dominated convergence (since $\int_{0}^{1}h(u)\,du<\infty$). Thus, for the given $t>0$ and $\delta\in(0,1]$, one can choose $M$ such that
$\frac{1}{p}\int_{0}^{p}h(u)\,\mathbf{1}_{\{h(u)>M\}}\,du\leq\frac{\delta t}{6},$
which implies both
$\mathbb{P}\Bigl(\bigl|\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]-b_{\mu^{(n)}}(p)\bigr|\geq t/3\Bigr)=0,$
and, by Markov’s inequality and (2.2),
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-\bar{X}^{(n,M)}\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)\leq\frac{3}{t}\,\mathbb{E}\bigl[|\bar{X}^{(n)}-\bar{X}^{(n,M)}|\mid\epsilon=1\bigr]\leq\frac{\delta}{2},$
uniformly in $n$.
Finally, on $\{\epsilon=1\}$ the truncated variables satisfy $X_{i}^{(n,M)}\in[-M,M]$ and remain independent; thus Hoeffding’s inequality gives
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n,M)}-\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)\leq 2\exp\!\Bigl(-\frac{nt^{2}}{18M^{2}}\Bigr)\leq\frac{\delta}{2},$ (14)
for $n$ large enough. Plugging these bounds into (12) yields
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-b_{\mu^{(n)}}(p)\bigr|\geq t\ \bigm|\ \epsilon=1\Bigr)\leq\delta,$
which proves $\bar{X}^{(n)}\mid\{\epsilon=1\}\to b_{\mu^{(n)}}(p)$ in probability. The case $\epsilon=0$ is identical and yields $\bar{X}^{(n)}\mid\{\epsilon=0\}\to a_{\mu^{(n)}}(p)$.
Step 4: Tail profile away from the discontinuity points. Let $\eta>0$ and $(t_{n})_{n}$ be as in the statement.
If $t_{n}\leq a_{\mu^{(n)}}(p)-\eta$, then on $\{\epsilon=0\}$ we have $\bar{X}^{(n)}\to a_{\mu^{(n)}}(p)$ in probability and therefore $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to 1$; similarly on $\{\epsilon=1\}$ since $b_{\mu^{(n)}}(p)\geq a_{\mu^{(n)}}(p)$, we also have $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)\to 1$. Hence $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 1$.
If $t_{n}\geq b_{\mu^{(n)}}(p)+\eta$, then $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)\to 0$ and $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to 0$, hence $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 0$.
Finally, if $a_{\mu^{(n)}}(p)+\eta\leq t_{n}\leq b_{\mu^{(n)}}(p)-\eta$, then $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)\to 1$ and $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to 0$, hence
$\mathbb{P}(\bar{X}^{(n)}\geq t_{n})=p\,\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)+(1-p)\,\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to p.$
This proves the three limits announced.